Given the following, calculate the Net Mortality Rate: In order to receive credit, the rate must be expressed as a percentage and be calculated to the hundreths place. In addition, you must show your work in order to receive credit.

Inpatient deaths: 52

Death Less than 48 hours after admission 13

Discharges (including deaths) 2076

Total autopsies 7

Bodies removed by medical examiner 1
Answer #1

ANSWER:

Given that,

Total Discharge= 2076 (including Death)

No. of Death = (52+13+7+1)=73

Number of death in 2076 People = 73

\therefore No. of death in 1 people = 73/ 2076

\therefore Morality rate =Number of death in 1000 people

= 73000/2076

= 35.16

%age Morality rate= 3.515%.